OpenStudy (anonymous):

Which statement describes the translation of y = – (x – 2)2 – 2 from standard position?

OpenStudy (paxpolaris):

translation by (h,k):\[y-k= f(x-h)\]

OpenStudy (anonymous):

i need this kind of answer Moved up and to the right. Moved up and to the left. Moved down and to the left. Moved down and to the right.

OpenStudy (paxpolaris):

right, so:\[f(x)=-x^2\] h=2, k =-2 So right by 2, down by 2

OpenStudy (anonymous):

thank you